Dana is 20 years older than daughter Kim. Ten years ago, Dana was twice as old as Kim was then. How old are Dana and Kim now?
Answer by chana(69)

Dana is 20 years older than daughter Kim. Ten years ago, Dana was twice as old as Kim was then. How old are Dana and Kim now?
sol:
let daughter age = x ,dana age = 20 + x
given :Ten years ago, Dana was twice as old as kim
(20 + x) - 10 = 2*(x - 10)
solving for x you get kim = 30 yrs & dana = 50 yrs.